OpenStudy (anonymous):

Please solve each of the inequalities below and graph the solution on a number line 1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

if you show work it will be such a great help

myininaya (myininaya):

Say we look at the last one together: \[-2(x-3) \ge 5-(x+3)\] And then you try the others... Ok... So do you know the distributive property a(b+c)=ab+ac

myininaya (myininaya):

Do you know how to apply that here: -2(x-3) ?